The wave length (W) of a radio wave varies inversely as its frequency (F). A wave length with a frequency of 1200 kilohertz has a length of 300 meters. What is the length of a wave with a frequency of 800 kilohertz?
Answer by stanbon(75887)

The wave length (W) of a radio wave varies inversely as its frequency (F). A wave length with a frequency of 1200 kilohertz has a length of 300 meters. What is the length of a wave with a frequency of 800 kilohertz?
--------------------
W = k/F
---------
Find "k":
300 = k/1200
k = 360000
--------
EQUATION:
W = 360000/F
---------
Find W if f = 800
W = 360000/800
W = 450 meters
===================
Cheers,
Stan H.